# MegaFusion: Extend Diffusion Models towards Higher-resolution Image Generation without Further Tuning

### v1

@@ Abstract
Diffusion models have emerged as frontrunners in text-to-image generation for their impressive capabilities. Nonetheless, their fixed image resolution during training often leads to challenges in high-resolution image generation, such as semantic inaccuracies and object replication. This paper introduces MegaFusion, a novel approach that extends existing diffusion-based text-to-image generation models towards efficient higher-resolution generation without additional fine-tuning or extra adaptation. Specifically, we employ an innovative truncate and relay strategy to bridge the denoising processes across different resolutions, allowing for high-resolution image generation in a coarse-to-fine manner. Moreover, by integrating dilated convolutions and noise re-scheduling, we further adapt the model’s priors for higher resolution. The versatility and efficacy of MegaFusion make it universally applicable to both latent-space and pixel-space diffusion models, along with other derivative models. Extensive experiments confirm that MegaFusion significantly boosts the capability of existing models to produce images of megapixels and various aspect ratios, while only requiring about $\mathbf{40\%}$ of the original computational cost. Code is available at https://haoningwu3639.github.io/MegaFusion/.
@@ 1 Introduction
Diffusion models have recently demonstrated unparalleled performance across broad applications, including text-to-image generation [15, 40, 16, 14, 35, 6], image editing [3, 28, 21, 12, 45, 26], consistent image sequence generation [25, 27, 29], and even achieves promising results in challenging text-to-video generation [17, 46, 14, 39]. Among them, the open-source Stable Diffusion (also known as Latent Diffusion [35]) leverages a pre-trained VAE [22] to compress images into a latent space for efficiency. After pre-training on large-scale paired datasets like LAION-5B [38], it has showcased impressive generative capabilities, thus being widely employed. In comparison, Imagen [37] and Floyd [6] opt for a cascading diffusion model in pixel space, initiating with low-resolution image synthesis followed by successive super-resolution stages to enable high-resolution text-to-image generation.
@@
Despite these advancements, as depicted in Figure 1 Left, these models face a significant hurdle: they struggle to generate images beyond the training range’s resolution, leading to semantic deviation and a decline in image quality. Existing efforts to address these issues often require extra tuning or are limited to specific models. For example, MultiDiffusion [2] and ElasticDiffusion [10] adopt post-processing or optimization techniques to stitch native-resolution images into higher-resolution panoramas, which is inefficient and time-consuming. Relay Diffusion [41] employs blurring diffusion and block noise in pixel space, yet it necessitates training multiple specific diffusion models from scratch. ResAdapter [5] and CheapScaling [9] introduce minimal extra training through LoRA [19] layers or Upsamplers, but still incur a non-negligible training overhead. Meanwhile, tuning-free alternatives like ScaleCrafter [11] and FouriScale [20] adapt pre-trained SDMs for higher resolution synthesis, but require meticulous hyperparameter adjustment and are restricted to latent-space models.
@@
To tackle the aforementioned limitations, we introduce MegaFusion, a powerful tuning-free method to extend existing diffusion models towards generating higher-resolution and variable aspect ratio images with megapixels. Concretely, we begin with a truncate and relay strategy, which seamlessly bridges the synthesis of different resolution images, enabling efficient higher-resolution generation in a coarse-to-fine manner with only about $40\%$ of the original computational cost. Additionally, we further improve the quality and details of generated images through dilated convolutions [49] and noise re-scheduling. The versatility and effectiveness of our MegaFusion make it compatible with both latent-space and pixel-space diffusion models, and its foundational principles can be extended to other diffusion-based frameworks with extra conditions, such as IP-Adapter [48] and ControlNet [51]. As shown in Figure 1 Right, MegaFusion significantly improves the ability of diffusion models to synthesize higher-resolution images that are both semantically accurate and of superior quality.
@@
To summarize, we make the following contributions in this paper: (i) we propose a tuning-free approach, MegaFusion, that utilizes a truncate and relay strategy to efficiently generate high-quality, high-resolution images with megapixels in a coarse-to-fine manner; (ii) we incorporate dilated convolution and noise re-scheduling techniques to further refine the adaptability of pre-trained diffusion models for higher resolution; (iii) we demonstrate the applicability of our method across both latent-space and pixel-space diffusion models, as well as their extensions, enabling the generation of high-resolution images with various aspect ratios at roughly 40% of the original computational expense; (iv) we conduct extensive experiments to validate the effectiveness and superiority of our proposed method, in terms of efficiency, image quality, and semantic accuracy.
@@ 2 Related Works
Diffusion Models. As a groundbreaking family in probabilistic generative models, diffusion models progressively inject noise into data in a forward process and learn to invert this process to generate new samples. DDPM [15] has showcased remarkable performance, outperforming previous GAN-based and auto-regressive techniques, while DDIM [40] significantly improves sampling efficiency. Leveraging their excellent generative capabilities, diffusion models have been applied to diverse fields, including image-to-image translation [3, 28, 21, 12, 45, 26], image sequence generation [25, 27, 29] and video generation [17, 46, 14, 39].
Text-to-Image Generation. Generative models have been widely adopted for the challenging text-to-image generation task, with GAN [8] and its variants [50, 47] as the pioneers. Meanwhile, auto-regressive transformers like DALL·E [34] further push the boundaries. Recently, diffusion models, such as DALL·E 2 [33], Imagen [37] and Floyd [6], have risen to prominence in text-to-image generation tasks. Notably, Stable Diffusion (Latent Diffusion Models [35]), performing the diffusion process in latent space, has demonstrated outstanding performance after extensive pre-training on vast image-text pair datasets, thereby being widely applied within the research community. Additionally, SDXL [30] further elevates the generative performance with a diffusion refiner and an extra text encoder.
@@
Higher-resolution Generation. Although diffusion models have shown promising performance in image generation, they are typically limited to fixed resolutions and aspect ratios, struggling to synthesize higher-resolution images beyond their training data. MultiDiffusion [2] addresses this by synthesizing overlapping crops with a pre-trained diffusion model and merging these to produce panoramic images, but it requires a time-consuming inference procedure. ElasticDiffusion [10] adopts a similar approach and improves image quality via reference images and reduced-resolution guidance. Relay Diffusion [41] constructs a pixel-space diffusion model with blurring diffusion and block noise to craft high-resolution images, at the cost of retraining several models from scratch. ScaleCrafter [11] improves high-resolution generation by addressing the UNet’s insufficient receptive field with dispersed convolution, without additional fine-tuning on Stable Diffusion. DemoFusion [7] attempts to connect multiple resolutions for coarse-to-fine generation, but it requires repeating multiple image generation processes, leading to low inference efficiency.
Several concurrent works also express rich interest in higher-resolution generation: ResAdapter [5] introduces a trainable adapter to generate images with unrestricted resolutions and aspect ratios by understanding resolution priors. CheapScaling [9] integrates multi-scale upsampler modules for multi-resolution generation with minimal tuning. FouriScale [20] and HiDiffusion [52] offer a training-free strategy but are still limited to SDM-based models.
In contrast to the aforementioned methods, which either necessitate further training or are limited to specific models, our proposed MegaFusion emerges as a versatile solution. It can be seamlessly integrated into both latent-space and pixel-space diffusion models, without any extra tuning or adaptation, synthesizing higher-resolution images of various aspect ratios with megapixels.
@@ 3.1 Diffusion Models
Diffusion models, a type of deep generative models, fundamentally transform input Gaussian noise into structured data samples through iterative denoising. Concretely, diffusion models comprise a forward diffusion process that progressively adds Gaussian noise to an image $\mathbf{x}_{0}$ via a Markov process over $T$ steps. Let $\mathbf{x}_{t}$ represent the noisy image at step $t$, with the transition from $\mathbf{x}_{t-1}$ to $\mathbf{x}_{t}$ being modeled by $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$. Here, $\beta_{t}\in(0,1)$ are pre-determined hyperparameters controlling the variance introduced at each step.
Defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, we can leverage the properties of Gaussian distributions and the reparameterization trick to reformulate the relationship as: $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$. This insight allows us to succinctly express the forward process with Gaussian noise $\epsilon$ as: $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$.
Diffusion models also encompass a reverse diffusion process to reconstruct an image from noise. This process, denoted as $p_{\theta}$, usually leverages a UNet-based [36] model to estimate the noise term $\epsilon_{\theta}$, represented as: $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$.
Here, $\mu_{\theta}$ represents the predicted mean value of Gaussian distribution, which can be expressed in terms of the noise $\epsilon_{\theta}$ as: $\mu_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t))$
@@ 3.2 Latent Diffusion Models
To improve efficiency and reduce computational cost, Latent Diffusion (LDMs) execute diffusion and denoising within a learned low-dimensional latent space of a pre-trained Variational Autoencoder (VAE). Specifically, the VAE encoder $\mathcal{E}$ converts an image $\mathbf{x}_{0}\in\mathbb{R}^{3\times H\times W}$ into a latent representation $\mathbf{z}_{0}\in\mathbb{R}^{4\times h\times w}$ via $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$. Afterwards, the decoder $\mathcal{D}$ reconstructs the original image $\mathbf{x}_{0}$ from $\mathbf{z}_{0}$, represented as $\hat{\mathbf{x}}_{0}=\mathcal{D}(\mathbf{z}_{0})\approx\mathbf{x}_{0}$.
This setup allows the diffusion process to be conducted in a compact latent space, enabling efficient image synthesis. During inference, LDM samples latent codes from a conditional distribution $p(\mathbf{z}_{0}|c)$, where $c$ represents the conditional information such as text embedding from CLIP [31] or T5 [32] text encoder. This process can be formalized as: $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},c)=\mathcal{N}(\mathbf{z}_{t};\mu_{\theta}(\mathbf{z}_{t},t,c),\Sigma_{\theta}(\mathbf{z}_{t},t,c))$
@@ 4.1 Truncate and Relay Strategy
For clarity, we will explain our method using latent-space diffusion models as an example. As for diffusion models in pixel space, our method can be applied more straightforwardly and conveniently.
With a pre-trained Stable Diffusion Model (SDM) equipped with a denoiser $\epsilon_{\theta}$, a low-resolution latent code $\mathbf{z}_{0}^{1}\in\mathbb{R}^{4\times h_{1}\times w_{1}}$ can be synthesized within $T$ denoising steps, then decoded into an image $\mathbf{x}_{0}^{1}\in\mathbb{R}^{3\times H_{1}\times W_{1}}$ by the VAE decoder $\mathcal{D}$. Conditioned on a text prompt $c_{T}$, our goal is to generate a higher-resolution image $\mathbf{x}_{0}^{k}\in\mathbb{R}^{3\times H_{k}\times W_{k}}$ alongside its corresponding latent code $\mathbf{z}_{0}^{k}\in\mathbb{R}^{4\times h_{k}\times w_{k}}$, by linking the generation processes across different resolutions over a total of $T$ steps, where $T=\sum_{i=1}^{k}T_{i}$.
Given that our methodology is inherently tuning-free, we focus on the inference stage herein. We begin with a low-resolution generation through $T_{1}$ steps denoising:
$\mathbf{z}_{t}=\frac{1}{\sqrt{\alpha_{t}}}$ $(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{z_{t}},t,c_{T}))+\sigma_{t}\epsilon,$ (1)
@@
where $\sigma_{t}$ are pre-calculated coefficients and $\epsilon$ denotes noise sampled from a standard Gaussian distribution. Subsequently, at step $t_{1}=T-T_{1}+1$, we truncate the generation process and compute the approximate clean latent code $\hat{\mathbf{z}}_{t_{1}}^{1}\in\mathbb{R}^{4\times h_{1}\times w_{1}}$, which serves as a pivotal element for multi-resolutions bridging:
@@
Then, the upsampled image $\hat{\mathbf{x}}_{t_{1}}^{2}$ is re-encoded into to latent code $\hat{\mathbf{z}}_{t_{1}}^{2}\in\mathbb{R}^{4\times h_{2}\times w_{2}}$ via the VAE encoder $\mathcal{E}$ and perturbed with noise at the current step $t_{1}$ to relay the generation:
@@
Table 1: Quantitative comparison. We compare our boosted models on higher-resolution image generation with representative latent-space and pixel-space diffusion models on the MS-COCO [24] dataset. Within each unit, we denote the best performance in RED and the second-best performance in $\mathrm{BLUE}$.

### v2

@@ Abstract
Diffusion models have emerged as frontrunners in text-to-image generation, however, their fixed image resolution during training often leads to challenges in high-resolution image generation, such as semantic deviations and object replication. This paper introduces MegaFusion, a novel approach that extends existing diffusion-based text-to-image generation models towards efficient higher-resolution generation without additional fine-tuning or extra adaptation. Specifically, we employ an innovative truncate and relay strategy to bridge the denoising processes across different resolutions, allowing for high-resolution image generation in a coarse-to-fine manner. Moreover, by integrating dilated convolutions and noise re-scheduling, we further adapt the model’s priors for higher resolution. The versatility and efficacy of MegaFusion make it universally applicable to both latent-space and pixel-space diffusion models, along with other derivative models. Extensive experiments confirm that MegaFusion significantly boosts the capability of existing models to produce images of megapixels and various aspect ratios, while only requiring about $\mathbf{40\%}$ of the original computational cost. Code is available at https://haoningwu3639.github.io/MegaFusion/.
@@ 1 Introduction
Diffusion models have recently demonstrated unparalleled performance across broad applications, including text-to-image generation [18, 42, 19, 17, 37, 8, 10], image editing [4, 30, 24, 15, 48, 28, 5], consistent image sequence generation [27, 29, 31], and even achieves promising results in challenging text-to-video generation [20, 49, 17, 41]. Among them, the open-source Stable Diffusion (also known as Latent Diffusion [37]) performs denoising in a compressed latent space, and has showcased impressive generative capabilities after pre-training on large-scale paired datasets [40]. In comparison, Imagen [39] and Floyd [8] opt for cascading diffusion models in pixel space, initiating with low-resolution image synthesis followed by successive super-resolution stages for text-to-image generation.
@@
Despite these advancements, as depicted in Figure 1 Left, these models face a major challenge: they struggle to generate images beyond training resolutions, leading to semantic deviation and degraded image quality. Existing solutions often require additional tuning or are limited to specific models. For example, MultiDiffusion [3] and ElasticDiffusion [13] adopt post-processing optimization to stitch high-resolution panoramas, which is inefficient and time-consuming. Relay Diffusion [43] employs blurring diffusion in pixel space, yet it necessitates training multiple specific diffusion models from scratch. ResAdapter [7] and CheapScaling [12] involve minimal extra training through LoRA [22] or Upsamplers, but still incur a notable training overhead. Meanwhile, tuning-free alternatives like ScaleCrafter [14] and FouriScale [23] adapt pre-trained SDMs for higher resolutions, but demand meticulous hyperparameter adjustment and are restricted to latent-space models.
@@
To tackle the aforementioned limitations, we introduce MegaFusion, a tuning-free method to extend existing diffusion models towards generating higher-resolution and variable aspect ratio images with megapixels. Concretely, we begin with a truncate and relay strategy, which seamlessly bridges the synthesis of different resolution images, enabling efficient generation in a coarse-to-fine manner with only $40\%$ of the original computational cost. Moreover, it is orthogonally compatible with existing techniques such as dilated convolutions [52] and noise re-scheduling for better image quality. The effectiveness and versatility of MegaFusion make it applicable to both latent-space and pixel-space diffusion models, as well as other diffusion-based frameworks with extra conditions, such as IP-Adapter [51] and ControlNet [54]. As shown in Figure 1 Right, MegaFusion significantly improves the ability of diffusion models to synthesize higher-resolution images with accurate semantics and superior quality.
@@
To summarize, our contributions are fourfold: (i) we propose MegaFusion, a tuning-free approach utilizing a truncate and relay strategy to efficiently generate high-quality, high-resolution images with megapixels in a coarse-to-fine manner; (ii) we incorporate dilated convolution and noise re-scheduling techniques to further refine the adaptability of pre-trained diffusion models for higher resolution; (iii) we demonstrate the applicability of our method across both latent-space and pixel-space diffusion models, as well as their extensions, synthesizing high-resolution images with various aspect ratios at roughly 40% of the original computational cost; (iv) we conduct extensive experiments validating the superiority of our proposed method, in terms of efficiency, image quality, and semantic accuracy.
@@ 2 Related Works
Diffusion Models. As a part of probabilistic generative models, diffusion models typically learn to generate new samples by iterative denoising. DDPM [18] has first showcased remarkable performance, while DDIM [42] significantly improves sampling efficiency. Leveraging their excellent generative capabilities, diffusion models have been applied to diverse fields, including image-to-image translation [4, 30, 24, 15, 48, 28], image sequence generation [27, 29, 31] and video generation [20, 49, 17, 41].
Text-to-Image Generation. Generative models have been widely adopted for the challenging text-to-image generation task, with GAN [11, 53, 50] as the pioneers. Meanwhile, auto-regressive transformers like DALL·E [36] further push the boundaries. Diffusion models, such as DALL·E 2 [35], Imagen [39] and Floyd [8], have recently risen to prominence. Notably, Stable Diffusion (Latent Diffusion [37]), performing denoising in latent space, has demonstrated outstanding performance, thereby being widely applied within the research community. Additionally, SDXL [32] further elevates the generative performance of Stable Diffusion with a diffusion refiner and an extra text encoder.
@@
Higher-resolution Generation. Existing diffusion models are typically limited to fixed resolutions and aspect ratios, struggling to produce images beyond their training resolutions. MultiDiffusion [3] and ElasticDiffusion [13] address this by synthesizing overlapping crops and merging them into panoramic images, which requires a time-consuming inference procedure. Relay Diffusion [43] designs a pixel-space model with blurring diffusion to craft high-resolution images, at the cost of retraining several models from scratch. ScaleCrafter [14] achieves high-resolution generation by enlarging receptive fields of Stable Diffusion with dispersed convolution without extra fine-tuning. DemoFusion [9] attempts to connect multiple resolutions for coarse-to-fine generation, but demands repeating generation multiple times, leading to low efficiency.
Several concurrent works also express rich interest in this task: ResAdapter [7] and CheapScaling [12] can generate images with unrestricted resolutions and aspect ratios with minimal tuning via trainable LoRA adapters or Upsamplers. FouriScale [23] and HiDiffusion [55] offer a training-free strategy but are still limited to SDM-based models.
In contrast to the aforementioned methods, which either require further training or are limited to specific models, our proposed MegaFusion emerges as a versatile tuning-free solution that can be integrated seamlessly into existing diffusion models, enabling the synthesis of higher-resolution images of various aspect ratios.
@@ 3.1 Diffusion Models
Diffusion models, a class of deep generative models, iteratively transform Gaussian noise into structured data samples through a denoising process. Concretely, diffusion models comprise a forward diffusion process that progressively adds Gaussian noise to an image $\mathbf{x}_{0}$ via a Markov process over $T$ steps. Let $\mathbf{x}_{t}$ represent the noisy image at step $t$, with the transition from $\mathbf{x}_{t-1}$ to $\mathbf{x}_{t}$ being modeled by $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$. Here, $\beta_{t}\in(0,1)$ are pre-determined hyperparameters controlling the variance introduced at each step. By defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, we can leverage the properties of Gaussian distributions and the reparameterization trick to reformulate the relationship as: $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$. This insight allows us to succinctly express the forward process with Gaussian noise $\epsilon$ as: $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$.
Diffusion models also encompass a reverse diffusion process to reconstruct images from noise. This process, denoted as $p_{\theta}$, usually leverages a UNet-based [38] model to estimate the noise term $\epsilon_{\theta}$, represented as: $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$. Here, $\mu_{\theta}$ is the predicted mean of Gaussian distribution, expressed in terms of the estimated noise $\epsilon_{\theta}$ as: $\mu_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t))$
@@ 3.2 Latent Diffusion Models
To improve efficiency and reduce computational cost, Latent Diffusion (LDMs) execute diffusion and denoising within a learned low-dimensional latent space of a pre-trained Variational Autoencoder (VAE). Specifically, the VAE encoder $\mathcal{E}$ maps an image $\mathbf{x}_{0}\in\mathbb{R}^{3\times H\times W}$ to a latent representation $\mathbf{z}_{0}\in\mathbb{R}^{4\times h\times w}$ via $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$. Afterwards, the decoder $\mathcal{D}$ reconstructs the original image $\mathbf{x}_{0}$ from $\mathbf{z}_{0}$, represented as $\hat{\mathbf{x}}_{0}=\mathcal{D}(\mathbf{z}_{0})\approx\mathbf{x}_{0}$.
This setup allows the diffusion process to be conducted in a compact latent space, facilitating efficient image synthesis. During inference, LDM samples latent codes from a conditional distribution $p(\mathbf{z}_{0}|c)$, where $c$ represents the conditional information such as text embedding from CLIP [33] or T5 [34] text encoder. This process can be formalized as: $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},c)=\mathcal{N}(\mathbf{z}_{t};\mu_{\theta}(\mathbf{z}_{t},t,c),\Sigma_{\theta}(\mathbf{z}_{t},t,c))$.
@@ 4.1 Truncate and Relay Strategy
High-level Idea. As evidenced by eDiff-I [1], diffusion models synthesize semantics during early denoising steps and texture details in later steps. Our intuition and insight here are that: we should perform early-stage denoising at original inference resolutions to guarantee accurate semantics, followed by truncating and relaying at higher resolutions to continue later-stage denoising to produce texture details. This strategy extends the higher-resolution generation capabilities of pre-trained models, enabling the synthesis of high-quality images with precise semantics at low computational costs, and supports various aspect ratios.
Problem Setting. For clarity, we focus on latent-space diffusion models as an example. As for pixel-space diffusion models, our method can be applied more straightforwardly and conveniently. Given that our strategy is inherently tuning-free, we focus on the inference stage herein. Using a pre-trained Latent Diffusion (LDM) with a denoiser $\epsilon_{\theta}$, a low-resolution latent code $\mathbf{z}_{0}^{1}\in\mathbb{R}^{4\times h_{1}\times w_{1}}$ can be synthesized within $T$ denoising steps conditioned on a text prompt $c_{T}$, and then decoded into an image $\mathbf{x}_{0}^{1}\in\mathbb{R}^{3\times H_{1}\times W_{1}}$ by the VAE decoder $\mathcal{D}$. Our goal is to generate a higher-resolution image $\mathbf{x}_{0}^{k}\in\mathbb{R}^{3\times H_{k}\times W_{k}}$ alongside its latent code $\mathbf{z}_{0}^{k}\in\mathbb{R}^{4\times h_{k}\times w_{k}}$, by linking generation processes across different resolutions over a total of $T$ steps, where $T=\sum_{i=1}^{k}T_{i}$.
Truncate. To guarantee accurate semantics, we begin with a low-resolution generation through $T_{1}$ steps denoising:
$\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}$ $(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{z}_{t},t,c_{T}))+\sigma_{t}\epsilon,$ (1)
@@
where $\sigma_{t}$ are pre-calculated coefficients and $\epsilon$ denotes noise sampled from a standard Gaussian distribution. Subsequently, at step $t_{1}=T-T_{1}+1$, we truncate the generation process and compute the approximate clean latent code $\hat{\mathbf{z}}_{t_{1}}^{1}\in\mathbb{R}^{4\times h_{1}\times w_{1}}$, which serves as a pivotal element for multi-resolutions bridging via:
@@
Relay. To further enhance higher-resolution texture details, the upsampled image $\hat{\mathbf{x}}_{t_{1}}^{2}$ is then re-encoded into to latent code $\hat{\mathbf{z}}_{t_{1}}^{2}\in\mathbb{R}^{4\times h_{2}\times w_{2}}$ via the VAE encoder $\mathcal{E}$ and perturbed with noise at the current step $t_{1}$ to relay the generation:
@@
Table 1: Quantitative comparison. We compare our boosted models on higher-resolution generation with representative latent-space and pixel-space diffusion models on MS-COCO [26] dataset. RED represents best performance, and BLUE denotes second best performance.